The planet Uranus has 13 more moons than Neputune, Saturn has 2 more than twice the number of moons of Neptune. The total number of moons for these planets is 47. How many moons does each planet have?
Answer by mananth(16946)

The planet Uranus has 13 more moons than Neputune, Saturn has 2 more than twice the number of moons of Neptune. The total number of moons for these planets is 47. How many moons does each planet have?
Let Neptune have x moons
uranus has x+13 moons
saturn has 2x+2
Total moons =47
..
x+x+13 +2x+2=47
4x+15=47
4x=32
x=8
..
Neptune has 8 moons
Uranus has 8+13 = 21 moons
Saturn has 2x+2 = 18 moons
..
CHECK
8+18+21=47